OpenStudy (anonymous):

Calculate the value of x that makes the expression true; round the answer to the nearest ten-thousandth. e^x – 5 = 8 Select one: a. 9.2564 b. 6.5942 c. 5.9642 d. 2.5649

OpenStudy (agreene):

is it \[e^x-1=8~~~~~~or~~~~~~~e^{x-1}=8\]

OpenStudy (agreene):

1 = 5

OpenStudy (anonymous):

e^x - 5 = 8 e^x = 13 x = Log[13] = 2.5649

OpenStudy (anonymous):

log[e^x ] = x